i am asked if y varies directly with x and y=10 when x=15, find y when x=6. i think the answer is 4 is that correct
Answer by Alan3354(69443)

i am asked if y varies directly with x and y=10 when x=15, find y when x=6. i think the answer is 4 is that correct
---------------
y = kx
10 = k*15
k = 10/15 = 2/3
----------------
y = (2/3)*6
y = 4
--------- It's correct
Or,
10:15::y:6
15y = 60
y = 4
--------